OpenStudy (anonymous):

how to simply a fraction with a fraction on bottom. example 1over1/2

OpenStudy (anonymous):

\[\frac{1}{\frac{1}{2}}=1\times \frac{2}{1}=1\times 2=2\]

OpenStudy (anonymous):

but if you are thinking you don't need to do even this, since there are evidently two "one halves" in one